How many milliliters of 1.90 M Pb(NO3)2 Pb(NO3)2 will react with 40.0 mL of 1.40 M KCl?
1
Expert's answer
2020-12-01T14:24:35-0500

Pb(NO3)2 + 2KCl=PbCl2+ 2 KNO3

n(KCl) = c*V=1,4*40*10-3= 0,056 mole

n(Pb(NO3)2)= 0,056/2=0,028 mole

V(Pb(NO3)2)=n/c=(0,028/1,90)*103=14,74 mL

Answer: 14,74 mL Pb(NO3)2